OpenStudy (linda3):

A circle has a radius of 8.5in . A sector of the circle has a central angle of pi/3 radians. Find the area of the sector.

OpenStudy (linda3):

@satellite73 can you please help me out a bit on this problem? thanks!

OpenStudy (lncognlto):

The formula for the area of a sector is\[A = \frac{ 1 }{ 2 }r ^{2}\theta\] Where theta equals the value of the central angle in radians. Replacing the values in this formula with those given, where r = 8.5 and theta = pi/3, we get:\[A = \frac{ 1 }{ 2 }(8.5)^{2}(\frac{ \Pi }{ 3 })\]

OpenStudy (lncognlto):

Calculating this gives us: A = ?

OpenStudy (linda3):

ummm.... so you are saying ....

OpenStudy (lncognlto):

LOL Which part of it didn't you get? xD

OpenStudy (linda3):

i'm sorry but i can't understand all the "A=fraq{(1)(2)" stuff can you use the equation buttons so that it can show it properly? Sorry for this confusion lol, but than you so very much for taking your time and helping me out.

OpenStudy (linda3):

thank*

OpenStudy (lncognlto):

No worries. LOL I actually did use the equation thingy, but apparently it isn't working again. So I will type it out best as I can: A = (1/2)(r)^2(theta) A = (1/2)(8.5)^2(pi/3)

OpenStudy (linda3):

so this means A= 1/2 x r to the power of 2 x theta and then plot in your numbers and get A= 1/2 x 8.5 to the power of 2 x pi/3

OpenStudy (lncognlto):

Yup, that's correct. ^-^

OpenStudy (linda3):

that being said the answer would be 37.8!

OpenStudy (lncognlto):

And that would be correct too! xD

OpenStudy (linda3):

THANK YOU so much!!!! :)

OpenStudy (lncognlto):

You're welcome! :D